Suppose two hosts A and B are directly connected. Length of the link is 10,000 km and the transmission rate is 10 Mbps. The propagation speed of the link in 0.5 * 108 m/s. Based on this information answer the following questions.

(i) Suppose A sends a file of 100 MB size. How long does it take to send the file? Assume the file is sent continuously.

 

(ii) Suppose the original file is broken up into 100 packets where size of each packet is 1MB. B sends an ACK for each packet and A cannot send a packet until the previous packet is acknowledged. Transmission time of an ACK packet is negligible. How long does it take to send the file?   


1
Expert's answer
2021-11-07T01:44:42-0400
  1. = (1 x 10 power 7 =10000000) / ( 0.5 x 10 power 8 =50000000)

=10000000/50000000 = 0.2sec

2 = (1 x 8 x 10 power 9)=(1 x 10 power 8)

= 134,217,728/ 100,000,000

= 1.34217728 sec